# Dicke superradiant enhancement of the heat current in circuit QED

### v1

@@ Abstract
Collective effects, such as Dicke superradiant emission, can enhance the performance of a quantum device. Here, we study the heat current flowing between a cold and a hot bath through an ensemble of $N$ qubits, which are collectively coupled to the thermal baths. We find a regime where the collective coupling leads to a quadratic scaling of the heat current with $N$ in a finite-size scenario. Conversely, when approaching the thermodynamic limit, we prove that the collective scenario exhibits a parametric enhancement over the non-collective case. We then consider the presence of a third uncontrolled parasitic bath, interacting locally with each qubit, that models unavoidable couplings to the external environment. Despite having a non-perturbative effect on the steady-state currents, we show that the collective enhancement is robust to such an addition. Finally, we discuss the feasibility of realizing such a Dicke heat valve with superconducting circuits. Our findings indicate that in a minimal realistic experimental setting with two superconducting qubits, the collective advantage offers an enhancement of approximately $10\%$ compared to the non-collective scenario.
@@ I Introduction
Quantum thermodynamics Kosloff (2013); Benenti et al. (2017); Vinjanampathy and Anders (2016); Goold et al. (2016); Binder et al. (2019); Lostaglio (2019) is the study of heat and work management in quantum systems. Within the recent blooming of quantum technologies, the critical aspect of energy management becomes increasingly crucialAuffèves (2022), due to the natural interest in building energy-efficient quantum technologies and limiting associated energy waste. Optimizing the heat and energy transport performances of quantum devices is, therefore, an appealing direction. In this context, it is particularly relevant to investigate quantum heat transport in superconducting circuit Quantum Electrodynamics (circuit QED) Blais et al. (2021) which is among the most promising platforms for quantum technologies and quantum computation Arute et al. (2019). The heat current flowing across circuit QED devices has been recently measured across various designs Fornieri and Giazotto (2017); Maillet et al. (2020); Ligato et al. (2022); Gubaydullin et al. (2022).
Two recent circuit QED experiments have investigated the heat transport of microwave photons scattering off a single qubit, realizing a photonic heat valve Ronzani et al. (2018); Senior et al. (2020). In these experiments, a single transmon qubit was capacitively coupled to two microwave resonators, each of them in contact with a resistance acting as a heat bath. A natural question is therefore whether the performance of such a device could be enhanced in the presence of multiple qubits interacting among each other. Indeed, transport of heat and energy are well known to be sensitive probes of collective and many-body effects Pekola and Karimi (2021); Fazio et al. (1998); Kane and Fisher (1996); Jezouin et al. (2013); Bertini et al. (2021); Larzul and Schirò (2023). A notable example where collective phenomena result in an enhanced emission and super-extensive scaling is provided by the Dicke model Dicke (1954), where an ensemble of $N$ atoms in an optical cavity collectively radiates with a superextensive intensity that scales as $N^{2}$, i.e. enhanced by a factor $N$ with respect to ordinary fluorescence, where atoms emit independently. In the Dicke model, the electrical dipoles of the atoms synchronize thanks to their collective coupling to the optical cavity modes, leading to an enhanced emission which has been dubbed “superradiance” Raimond et al. (2001); Gross and Haroche (1982). Superradiant emission has been observed in various systems, including Rydberg atoms in cavities Kaluzny et al. (1983), color centers in diamonds Angerer et al. (2018), and in superconducting qubits Wang et al. (2020).
Collective effects, including Dicke superradiance, have been proposed to improve the performance of thermometers Correa et al. (2015); Mehboudi et al. (2022); Abiuso et al. (2023), of quantum heat engines Campisi and Fazio (2016); Fusco et al. (2016); Niedenzu and Kurizki (2018); Abiuso and Perarnau-Llobet (2020); Liu et al. (2023); Kamimura et al. (2023, 2023); Carrega et al. (2023); Tajima and Funo (2021), of quantum batteries Ferraro et al. (2018); Andolina et al. (2019); Quach et al. (2022); Campaioli et al. (2023, 2017), and to reduce the dissipated work Rolandi et al. (2023).
@@
Furthermore, we show that this collective enhancement is robust to the addition of a third uncontrolled parasitic bath, which interacts locally with each qubit. This finding is crucial since, as we will show, the presence of an infinitesimally small local noise has a finite and non-perturbative effect on the steady-state heat currents, and removes the dependence of the steady-state heat current from the initial state preparation. Furthermore, it demonstrates the resilience of the superradiant effect in a more realistic noisy environment.
@@
Finally, we discuss the feasibility of measuring such a Dicke enhancement of the heat current in an experimental setup. Such a device can be realized within the framework of circuit-QED Ronzani et al. (2018); Pekola (2015); Blais et al. (2021); Krantz et al. (2019), where, $N$ transmons are capacitively coupled to an ${\rm RLC}$ circuit where the dissipative nature of a thermal bath stems from the presence of the resistive elements. Our findings indicate that, in the minimal case of $N=2$ superconducting qubits, the collective advantage with experimentally realistic parameters is approximately $10\%$ higher than in scenarios where collective effects are absent. This highlights the potential of leveraging collective quantum behaviors in practical thermodynamic applications, offering a measurable enhancement over more traditional designs, and providing a platform to experimentally observe superradiant effects in measurable steady-steady heat currents.
Our proposed device also represents a many-body collective version of a heat valve, a device that has been previously implemented in the context of circuit-QED Ronzani et al. (2018). A heat valve is a device designed to control the flow of heat between two baths. Indeed, the circuit pictured in Fig. 2 can function as a heat valve. The ${\rm LC}$ circuits in Fig. 2 act as a filter, allowing the coupling only with frequencies near their resonant frequency $\omega_{\rm LC}$. This selective coupling is integral to the operation of our device as a heat valve. When the qubits are tuned to resonate at the $LC$ frequency, there is an efficient flow of heat between the two baths. Conversely, if the qubits are detuned from the $LC$ frequency, the heat flow is effectively impeded due to the frequency mismatch. This dynamic tuning capability allows for controlled manipulation of heat transfer, embodying the core function of a heat valve.
@@
This article is organized as follows, in Sec. II we derive the model and discuss the super-extensive behavior of the heat current in the noiseless case, i.e. when the modulus collective spin operator is conserved. In Sec. III we discuss the resilience of this super-extensive behavior to the addition of a realistic parasitic bath, breaking the conservation of the modulus of the collective spin. In Sec. V we draw our conclusions. Apps. A,B,C contain a series of technical details.
@@ II The Dicke-superradiant heat valve
The lumped-element circuit diagram, including capacitances, inductances, resistances, and the various variables, is depicted in Fig. 2. The quantum heat valve under study consists of $N$ transmon qubits coupled to two resistors denoted by $R_{i}$ in parallel with $LC$-resonators. These $RLC$ circuits form two heat baths characterized by their temperature $T_{i}$. We denote the two baths as ‘hot’ (h) and ‘cold’ (c), hence $i=\{\rm{h},\rm{c}\}$ and $T_{h}\geq T_{c}$. The total Hamiltonian of the system can be written as
$\hat{H}_{\rm tot}=\hat{H}_{0}+\hat{H}_{\rm int}+\hat{H}_{\rm RLC}\,.$ (1)
The first term describes the Hamiltonian of the qubits
$\hat{H}_{0}=\sum_{j=1}^{N}\frac{\hbar\omega_{0}}{2}\hat{\sigma}^{(j)}_{z}$ (2)
where $\hat{\sigma}^{(j)}_{z}$ represents the Pauli $z$-operator for the $j$-th qubit and $\hbar\omega_{0}$ is the qubit energy. The second term $\hat{H}_{\rm int}$ accounts for the coupling between the dipole operator of the $j$-th qubit $\hat{\sigma}^{(j)}_{x}$ and the voltage $\hat{V}_{i}$ of the $RLC$ circuit, and reads
$\hat{H}_{\rm int}=-\sum_{j=1}^{N}\sum_{i={\rm h,c}}{\hbar G_{i}}\frac{\hat{\sigma}^{(j)}_{x}}{2}\hat{V}_{i}$ (3)
where $G_{i}$ is the coupling strength that we consider different for the two RLC circuits but otherwise uniform across the qubits. The specific relationship of $G_{i}$ to the circuit’s parameters is detailed in Appendix A. This coupling plays a crucial role in the interaction dynamics of the qubits with their respective environments. Finally, $\hat{H}_{\rm RLC}$ describes the Hamiltonian of the two $RLC$ circuits as described in Appendix A, where we show that the total Hamiltonian of the system can be derived from a circuit QED setup. As shown in Appendix A, the Hamiltonian of the system can be derived from a circuit QED setup (see Eq. (58)), drawn in Fig. 2 and reads:
$\begin{split}\hat{H}_{0}&=\frac{1}{2}\sum_{j=1}^{N}\left(\frac{1}{C_{\rm eff}}\hat{Q}_{j}^{2}+\frac{1}{L_{\text{J}}}\hat{\phi}^{2}_{j}\right)~{},\\ \hat{H}_{\rm int}&=-\sum_{j=1}^{N}\sum_{i={\rm h,c}}\frac{C^{(i)}_{\rm c}}{C_{\rm T}}\hat{Q}_{j}\hat{V}_{i}~{},\end{split}$ (4)
where $C_{\rm eff}=C_{\rm T}\left[1-{\sum_{i={\rm h,c}}C^{(i)}_{\rm c}}/{C_{\rm T}}\right]^{-1}$, $\hat{Q}_{j},\hat{\phi}_{j}$ are the charge and flux associated with the $j$-th transmon, $C_{\rm T}$ ($L_{\rm J}$) is the transmon capacitance (inductance) and $C^{(i)}_{\rm c}$ are the coupling inductances, as reported in the lumped-circuit diagram in Fig. 2. By projecting the Hilbert space on the two-level subspace of each transmon we obtain the Hamiltonian in Eqs. (2,3).
@@
$\begin{split}\hat{H}_{\rm tot}=\hbar\omega_{0}\hat{J}_{z}-\sum_{i={\rm h,c}}{\hbar G_{i}}\hat{J}_{x}\hat{V}_{i}+\hat{H}_{\rm RLC}\,.\end{split}$ (6)
@@
where $J(J+1)$ ($m_{J}$) are the eigenvalues of $\hat{J}^{2}$ ($\hat{J}_{z}$), and $\hat{J}_{+}$ is the raising operator.
@@
We are interested in a scenario where the qubits are weakly coupled to the baths. In this regime, it is possible to invoke the Born-Markov approximation and describe the open system dynamics of the qubits with a suitable Lindblad master equation
@@
where $\hat{\rho}$ is the density operator of the qubits. The linear superoperator $\mathcal{D}(\hat{\rho})$, known as the dissipator, arises from integrating out the RLC circuits which are coupled to our system through the voltage $\hat{V}_{i}$. Within the Born-Markov approximation, the dynamics of the voltage is fully encoded in the voltage dynamical structure factor $S_{\hat{V}_{i},\hat{V}_{j}}(\omega)$ Cattaneo and Paraoanu (2021).
A detailed derivation of the Lindblad master equation is carried out in Appendix B, while here we summarize the results needed for our analysis. The interaction term in Eq. (6), coupling the qubits system with the two resonators, is proportional to $\hat{J}_{x}=\hat{J}_{+}+\hat{J}_{-}$. Hence, after performing the secular approximation needed to derive a Lindblad master equation Breuer and Petruccione (2007), the suitable Lindblad operators for our system, describing transitions between the Dicke states, can be conveniently written using the collective spin raising and lowering operators $\hat{J}_{+}$ and $\hat{J}_{-}$. Hence, the total dissipator in Eq. (9) can be separated into contributions from the hot and cold baths:
@@
where $i={\rm h,c}$ and $\gamma_{\rm c}$ and $\gamma_{\rm h}$ are the transition rates for the cold and hot baths, respectively. The rates $\gamma_{i}$ can be expressed in terms of the microscopic parameters of the circuit, as we show in Appendix  B. The thermal occupation numbers $n_{i}$ are given by the Bose–Einstein distribution
@@
To summarize, after integrating out the two RLC circuits representing the two thermal baths, we have obtained an effective description where the system of $N$ qubits is in contact with two structured thermal baths at finite temperature, each providing injection and depletion of energy through global spin-flip processes as described by Eq. (11). It is important to note that the Lindblad master equation in Eq. (9) correctly inherits the conservation of $\hat{J}^{2}$ from the total Hamiltonian in Eq. (6). Specifically, this becomes a strong symmetry of the Lindbladian superoperator, since $\hat{J}^{2}$ commutes both with $H_{0}$ as well as with each jump operator $\hat{J}_{\pm}$. As such this corresponds to a conserved quantity of the Lindblad dynamics Albert and Jiang (2014). Finally, we note that related Linbdlad master equations appeared in the description of dissipative (boundary) time crystals Iemini et al. (2018), where usually one considers a single bath at zero temperature.

### v3

# Dicke superradiant heat current enhancement in circuit quantum electrodynamics

@@ Abstract
Controlling the flow of energy and heat at the micro-scale is crucial to achieve energy-efficient quantum technologies, for on-chip thermal management, and to realize quantum heat engines and refrigerators. Yet, the efficiency of current quantum technologies is affected by thermal noise, and efficient cooling of quantum devices remains challenging in various solid-state implementations such as superconducting circuits. Collective effects, such as the Dicke superradiant emission, have been exploited to enhance the performance of quantum devices. However, the inherently transient nature of Dicke superradiant emission raises questions about its impact on steady-state properties. Here, we study how to enhance the steady-state heat current flowing between a hot and a cold bath through an ensemble of $N$ qubits, that are collectively coupled to the thermal baths. Remarkably, we find a regime where the heat current scales quadratically with $N$ in a finite-size scenario. Conversely, when approaching the thermodynamic limit, we prove that the collective scenario exhibits a parametric enhancement over the non-collective case. We then consider the presence of a third uncontrolled parasitic bath, interacting locally with each qubit, that models unavoidable couplings to the external environment. Despite having a non-perturbative effect on the steady-state currents, we show that the collective enhancement is robust to such an addition. Finally, we discuss the feasibility of realizing such a Dicke heat valve with superconducting circuits. Our findings indicate that in a minimal realistic experimental setting with two superconducting qubits, the collective advantage offers an enhancement of approximately $10\%$ compared to the non-collective scenario.
@@ I Introduction
Quantum thermodynamics [1, 2, 3, 4, 5, 6] is the study of heat and work management in quantum systems. Within the recent blooming of quantum technologies, the critical aspect of energy management becomes increasingly crucial[7], due to the natural interest in building energy-efficient quantum technologies and limiting associated energy waste. In this context, it is particularly relevant to investigate quantum heat transport in superconducting circuit Quantum Electrodynamics (circuit QED) [8] which is among the most promising platforms for quantum technologies and quantum computation [9]. The heat current flowing across circuit QED devices has been recently measured across various designs [10, 11, 12, 13, 14], and circuit QED is emerging as a platform to realize quantum heat engines and refrigerators [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Two recent circuit QED experiments have investigated the heat transport of microwave photons scattering off a single qubit, realizing a photonic heat valve [28, 29]. In these experiments, a single transmon qubit was capacitively coupled to two microwave resonators, each of them in contact with a resistance acting as a heat bath. A natural question is therefore whether the performance of such a device could be enhanced in the presence of multiple qubits interacting among each other. Indeed, transport of heat and energy are well known to be sensitive probes of collective and many-body effects [30, 31, 32, 33, 34, 35]. A notable example where collective phenomena result in an enhanced emission and super-extensive scaling is provided by the Dicke model [36], where an ensemble of $N$ atoms in an optical cavity collectively radiates with a superextensive intensity that scales as $N^{2}$, i.e. enhanced by a factor $N$ with respect to ordinary fluorescence, where atoms emit independently. In the Dicke model, the electrical dipoles of the atoms synchronize thanks to their collective coupling to the optical cavity modes, leading to an enhanced emission which has been dubbed “superradiance” [37, 38]. Superradiant emission has been observed in various systems, including Rydberg atoms in cavities [39], color centers in diamonds [40], and in superconducting qubits [41].
Collective effects, including Dicke superradiance, have been proposed to improve the performance of thermometers [42, 43, 44], of quantum heat engines [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 54, 55, 56], of quantum batteries [57, 58, 59, 60, 61], of refrigerators [62], of heat transfer[63], of energy transfer between quantum devices [64] and to reduce the dissipated work in finite-time thermodynamics [65].
@@
While many previous proposals have examined the influence of collective effects on thermal machines [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 64] and heat transport [63], it remains crucial to discuss the robustness of these effects to noise. Noise can hinder the coherence of the dynamics, potentially undermining the collective enhancement. Here, we show that this collective enhancement is robust to the addition of a third uncontrolled parasitic bath, which interacts locally with each qubit. This finding is crucial since, as we will show, the presence of an infinitesimally small local noise has a finite and non-perturbative effect on the steady-state heat currents, and removes the dependence of the steady-state heat current from the initial state preparation. Furthermore, we demonstrate the resilience of the superradiant effect in this realistic noisy environment.
@@
Despite the intense theoretical interest in exploiting collective effects in quantum thermodynamical machine [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 54, 55, 56, 57, 58, 60, 61, 62, 64, 63], only a few experiments have been conducted in this context [66, 59]. Hence, we discuss the feasibility of measuring such a Dicke enhancement of the heat current in an experimental setup. Such a device can be realized within the framework of circuit-QED [28, 67, 8, 68], where $N$ transmons are capacitively coupled to two $RLC$ circuits where the dissipative nature of a thermal bath stems from the presence of the resistive elements. Our findings indicate that, using experimentally realistic parameters in the minimal case of $N=2$ superconducting qubits, the collectively enhanced heat current is approximately $10\%$ higher than in scenarios where collective effects are absent. This highlights the potential of leveraging collective quantum behaviors in practical thermodynamic applications, offering a measurable enhancement over more traditional designs, and providing a platform to experimentally observe superradiant effects in measurable steady-state heat currents. Our proposed device also represents a many-body collective version of a heat valve, a device that has been previously implemented in the context of circuit-QED [28]. A heat valve is a device designed to control the flow of heat between two baths. Indeed, the circuit pictured in Fig. 2 can function as a heat valve. The ${\rm LC}$ circuits in Fig. 2 (left and right elements) act as a filter, when the qubits are tuned to resonate at the $LC$ frequency, there is an efficient flow of heat between the two baths. Conversely, if the qubits are detuned from the $LC$ frequency, the heat flow is effectively impeded due to the frequency mismatch. This dynamic tuning capability allows for controlled manipulation of heat transfer, embodying the core function of a heat valve.
@@
This article is organized as follows, in Sec. II we derive the model and discuss the super-extensive behavior of the heat current in the noiseless case, i.e. when the modulus of the collective spin operator is conserved. In Sec. III we discuss the resilience of this super-extensive behavior to the addition of a realistic parasitic bath, breaking the conservation of the modulus of the collective spin. In Sec. IV we discuss the experimental feasibility of observing our findings and we propose a minimal experimental setup that exhibits the super-extensive behavior of the heat current. In Sec. V we draw our conclusions. Apps. A,B,C, D contain a series of technical details.
@@ II The Dicke-superradiant heat valve
The lumped-element circuit diagram, including capacitances, inductances, resistances, and the various variables, is depicted in Fig. 2. The quantum heat valve under study consists of $N$ transmon qubits coupled to two resistors denoted by $R_{i}$ in parallel with $LC$-resonators. These $RLC$ circuits form two heat baths characterized by their temperature $T_{i}$. We denote the two baths as ‘hot’ (h) and ‘cold’ (c), hence $i=\{\rm{h},\rm{c}\}$ and $T_{h}\geq T_{c}$. As derived in App. A, the total Hamiltonian of the system can be written as
$\hat{H}_{\rm tot}=\hat{H}_{0}+\hat{H}_{\rm int}+\hat{H}_{RLC}\,.$ (1)
$\hat{H}_{0}$ and $\hat{H}_{\rm int}$ describe, respectively, the Hamiltonian of the qubits and their coupling to the $RLC$ circuit. Before projecting onto the qubit subspace, the Hamiltonian of the circuit is given by
$\begin{split}\hat{H}_{0}&=\frac{1}{2}\sum_{j=1}^{N}\left[\frac{1}{C_{\rm eff}}\hat{Q}_{j}^{2}-E_{\text{J}}\cos\left(\frac{2\pi\hat{\phi}_{j}}{\phi_{0}}\right)\right]~{},\\ \hat{H}_{\rm int}&=-\sum_{j=1}^{N}\sum_{i={\rm h,c}}\frac{C^{(i)}_{\rm c}}{C_{\rm T}}\hat{Q}_{j}\hat{V}_{i}~{},\end{split}$ (2)
where $C_{\rm eff}=C_{\rm T}\left[1-{\sum_{i={\rm h,c}}C^{(i)}_{\rm c}}/{C_{\rm T}}\right]^{-1}$, $\hat{Q}_{j},\hat{\phi}_{j}$ are the charge and flux associated with the $j$-th transmon, $C_{\rm T}$ ($E_{\text{J}}$) is the transmon capacitance (Josephson energy), $\phi_{0}=h/(2e)$ , $C^{(i)}_{\rm c}$ are the coupling inductances, and $\hat{V}_{i}$ is the voltage of the $i$-th $RLC$ circuit (for $i=\text{h},\text{c}$), as reported in the lumped-circuit diagram in Fig. 2. By projecting the Hilbert space onto the two-level subspace of each transmon we obtain
$\hat{H}_{0}=\sum_{j=1}^{N}\frac{\hbar\omega_{0}}{2}\hat{\sigma}^{(j)}_{z},$ (3)
where $\hat{\sigma}^{(j)}_{z}$ represents the Pauli $z$-operator for the $j$-th qubit and $\hbar\omega_{0}$ is the qubit energy, and
$\hat{H}_{\rm int}=-\sum_{j=1}^{N}\sum_{i={\rm h,c}}{\hbar G_{i}}\frac{\hat{\sigma}^{(j)}_{x}}{2}\hat{V}_{i},$ (4)
where $\hat{\sigma}^{(j)}_{x}$ is the dipole operator of the $j$-th qubit (the Pauli $x$-operator) and $G_{i}$ is the coupling strength that we consider different for the two $RLC$ circuits but otherwise uniform across the qubits. This coupling plays a crucial role in the interaction dynamics of the qubits with the thermal baths. Finally, $\hat{H}_{RLC}$ describes the Hamiltonian of the two $RLC$ circuits as described in Appendix A.
@@
$\begin{split}\hat{H}_{\rm tot}=\hbar\omega_{0}\hat{J}_{z}-\sum_{i={\rm h,c}}{\hbar G_{i}}\hat{J}_{x}\hat{V}_{i}+\hat{H}_{RLC}\,.\end{split}$ (6)
@@
where $J(J+1)$ ($m_{J}$) are the eigenvalues of $\hat{J}^{2}$ ($\hat{J}_{z}$), and $\hat{J}_{+}$ ($\hat{J}_{-}$) is the raising (lowering) operator.
@@
Motivated by existing experimental setups [28], we are interested in a scenario where the qubits are weakly coupled to the baths. In this regime, it is possible to invoke the Born-Markov approximation and describe the open system dynamics of the qubits with a suitable Lindblad master equation
@@
where $\hat{\rho}$ is the density operator of the qubits. A detailed derivation of this Lindblad master equation is carried out in Appendix B, while here we summarize the results needed for our analysis. The dissipator $\mathcal{D}(\hat{\rho})$ arises from integrating out the $RLC$ circuits which are coupled to our system through the voltage $\hat{V}_{i}$ (see Eqs. (4) and (6)). Within the Born-Markov approximation weak-coupling approximation, the dynamics of the voltage is fully encoded in the voltage dynamical structure factor [69] $S_{\hat{V}_{i},\hat{V}_{j}}(\omega)$, defined in Eq. (65), which describe the voltage fluctuations occurring in the $RLC$ circuits.
The interaction term in Eq. (6), coupling the qubits system with the two resonators, is proportional to $\hat{J}_{x}=\hat{J}_{+}+\hat{J}_{-}$. Hence, after performing the secular approximation as needed to derive a Lindblad master equation [70], the suitable Lindblad operators for our system, describing transitions between the Dicke states induced by the exchange of energy with thermal baths, can be conveniently written using the collective spin raising and lowering operators $\hat{J}_{+}$ and $\hat{J}_{-}$. Hence, the total dissipator in Eq. (9) can be separated into contributions from the hot and cold baths:
@@
where $i={\rm h,c}$ and $\gamma_{\rm c}$ and $\gamma_{\rm h}$ are the transition rates for the cold and hot baths, respectively. The rates $\gamma_{i}$, that represent the coupling strength between the baths and the qubits, can be expressed in terms of the microscopic parameters of the circuit, as shown in Appendix B. The thermal occupation numbers $n_{i}$ are given by the Bose–Einstein distribution
@@
To summarize, after integrating out the two $RLC$ circuits representing the two thermal baths, we have obtained an effective description where the system of $N$ qubits is in contact with two structured thermal baths at finite temperature, each providing injection and depletion of energy through global spin-flip processes as described by Eq. (11). It is important to notice that the Lindblad master equation in Eq. (9) correctly inherits the conservation of $\hat{J}^{2}$ from the total Hamiltonian in Eq. (6). Specifically, this becomes a strong symmetry of the Lindbladian superoperator, since $\hat{J}^{2}$ commutes both with $H_{0}$ as well as with each jump operator $\hat{J}_{\pm}$. As such this corresponds to a conserved quantity of the Lindblad dynamics [71]. Interstingly, analogous Lindblad master equations have been derived to describe dissipative (boundary) time crystals [72], where usually one considers a single bath at zero temperature.